OpenStudy (anonymous):

when doing this problem how do you simplify? 6k-2=2;k=2/3

OpenStudy (bahrom7893):

6k - 2 = 2 Add 2 to both sides: 6k - 2 + 2 = 2 + 2 6k = 4 k = 4/6, divide both top and bottom by 2 to get: k = 2/3